The perimeter of a rectangle is30m. If lenght exceed its breadth by 3m then find its lenght
Answer by Herdaysoji(25)

Perimeter of rec =l+l+b+b or 2(l+b)
.'. 30=2(l+b)
since l =b+3
30=2(3+b+b)
30=2(3+2b)
30=6+4b
4b=30-6
4b=24
b=24/4
b=6 and
l=3+6
l=9